OpenStudy (anonymous):

Will medal & Fan. Determine the solution set of x^2 - 80 = 0. a.) +or- 2 sqrt 5 b.) +or- sqrt 5 c.) +or- 4 sqrt 5

OpenStudy (hybrik):

Move 80 to other side

OpenStudy (hybrik):

x^2=80

OpenStudy (anonymous):

okay

OpenStudy (hybrik):

Then simplify the root of 80

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4 * 20?

OpenStudy (hybrik):

\[\sqrt{80} = \sqrt{20*4} = 2\sqrt{20} = 2\sqrt{4*5} = 2*2\sqrt{5} = 4\sqrt{5}\]

OpenStudy (anonymous):

oh thanks :)

OpenStudy (anonymous):

√=20∗4−−−−−√=220−−√=24∗5−−−−√=2∗25√=45√

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

ill fan m8. gave to hyb already.

OpenStudy (hybrik):

Juan, you kind of did it wrong?